Sales Budget

FlashKick Company manufactures and sells soccer balls for teams of children in elementary and high school. FlashKick’s best-selling lines are the practice ball line (durable soccer balls for training and practice) and the match ball line (high-performance soccer balls used in games). In the first four months of next year, FlashKick expects to sell the following:

Practice Balls Match Balls
Units Selling Price Units Selling Price
January 50,000 $8.25 7,000 $15.00
February 56,000 $8.25 7,500 $15.00
March 80,000 $8.25 13,000 $15.00
April 100,000 $8.25 18,000 $15.00

Required:

1. Construct a sales budget for FlashKick for the first three months of the coming year. Show total sales for each product line by month and in total for the first quarter. If required, round your answers to the nearest cent.

FlashKick Company
Sales Budget
For the First Quarter of Next Year
January February March Quarter
Practice ball:
Units
Unit price $ $ $ $
Sales $ $ $ $
Match ball:
Units
Unit price $ $ $ $
Sales $ $ $ $
Total sales $ $ $ $

2. What if FlashKick added a third line—tournament quality soccer balls that were expected to take 40 percent of the units sold of the match balls and would have a selling price of $42 each in January and February, and $45 each in March? Prepare a sales budget for FlashKick for the first three months of the coming year. Show total sales for each product line by month and in total for the first quarter. If required, round your answers to the nearest cent.

FlashKick Company
Sales Budget
For the First Quarter
January February March Quarter
Practice ball:
Units
Unit price $ $ $ $
Sales $ $ $ $
Match ball:
Units
Unit price $ $ $ $
Sales $ $ $ $
Tournament ball:
Units
Unit price $ $ $ $
Sales $ $ $ $
Total sales $ $ $ $
Answer #1
1)
FlashKick Company
Sales Budget
For the First Quarter of Next Year
January (x) February (y) March (z) Quarter (x + y + z)
Practice ball:
Units (a) 50,000 56,000 80,000 186,000
Units Price (b) $8.25 $8.25 $8.25 $8.25
Sales (c = a * b) $412,500 $462,000 $660,000 $1,534,500
Match ball:
Units (d) 7,000 7,500 13,000 27,500
Units Price (e ) $15 $15 $15 $15
Sales (f = d * e) $105,000 $112,500 $195,000 $412,500
Total Sales (c + f) $517,500 $574,500 $855,000 $1,947,000
2)
FlashKick Company
Sales Budget
For the First Quarter of Next Year
January (x) February (y) March (z) Quarter (x + y + z)
Practice ball:
Units (a) 50,000 56,000 80,000 186,000
Units Price (b) $8.25 $8.25 $8.25 $8.25
Sales (c = a * b) $412,500 $462,000 $660,000 $1,534,500
Match ball:
Units (d) 7,000 7,500 13,000 27,500
Units Price (e ) $15 $15 $15 $15
Sales (f = d * e) $105,000 $112,500 $195,000 $412,500
Tournament ball:
Units (g) (7,000*40/100); (7,500*40/100); (13,000*40/100); 2,800 3,000 5,200 11,000
Units Price (h) $42 $42 $45 $43.42
Sales (i = g * h) $117,600 $126,000 $234,000 $477,600
Total Sales (c + f + i) $635,100 $700,500 $1,089,000 $2,424,600